OpenStudy (dan815):

heelp for the surface with parametric equations r(s,t)=, find the equation of the tangent plane at (2,3,1) Also Find the surface area under the restriction s^2+t^2<=1.

OpenStudy (goformit100):

heelp for the surface with parametric equations r(s,t)=<st, s+t, s-t>, find the equation of the tangent plane at (2,3,1) Also Find the surface area under the restriction s^2+t^2<=1.

OpenStudy (anonymous):

do you know how to find the normal vector of a parametric surface?

OpenStudy (anonymous):

Do you know how to find a plane given a normal vector and a point?

OpenStudy (dan815):

no i only know if its in x y z

OpenStudy (anonymous):

Give me a second... brb

OpenStudy (anonymous):

Okay, the normal vector to the parametric surface \(\mathbf{r}(s,t)\) Is given by \[\large \mathbf{r}_s\times \mathbf{r}_t \]That is, the cross product each of its partial derivatives.

OpenStudy (anonymous):

@dan815 Think you can do that part at least?

OpenStudy (anonymous):

You will get the normal vector as a function of \(s,t\)

OpenStudy (anonymous):

Since the normal vector is changing

OpenStudy (dan815):

-2, t+s, t-s

OpenStudy (dan815):

wut do i do after i get the nomal vector

OpenStudy (anonymous):

Find \(s,t\) such that \(\mathbf{r}(s,t)=(2,3,1) \)

OpenStudy (anonymous):

Then plug in that \(s,t\) to find the normal vector to the point \((2,3,1) \)

OpenStudy (dan815):

ohh ok thanks i get it now

OpenStudy (dan815):

-2x+3y-z=4